# Invertibility of the Fourier Diffraction Relation in Raster Scan Diffraction Tomography

### v1

@@ Abstract
Diffraction tomography aims to recover an object’s scattering potential from measured wave fields. In the classical setting, the object is illuminated by plane waves from many directions, and the Fourier diffraction theorem gives a direct relation between the Fourier transform of the scattering potential of the object and the Fourier transformed measurements.
In many practical imaging systems, however, focused beams are used instead of plane waves. These beams are then translated across the object to bring different regions of interest into focus. The Fourier diffraction relation adapted to this setting differs in one crucial point from the plane-wave case: while certain Fourier coefficients of the measurements still directly correspond to individual Fourier coefficients of the scattering potential, others are given by linear combinations of two Fourier coefficients of the scattering potential.
This article investigates which Fourier coefficients of the scattering potential can be uniquely recovered from these relations. We show that in dimensions higher than two, all coefficients appearing in the equations are typically uniquely determined. In two dimensions, however, only part of the Fourier coverage is uniquely recoverable, while on the remaining subset distinct coefficients can produce identical data.
@@ 1. Introduction
To obtain the involved Fourier data of the scattering potential, we therefore still have to solve a linear equation system. But before doing so, we want to check in this article for every scan setup which Fourier coefficients are in fact uniquely determined by these equations.
While the physically relevant case is cleary the three-dimensional setting, we would like to discuss the problem in arbitrary dimensions, with particular attention to the theoretically interesting two-dimensional case where the dimensions of the gathered data and the Fourier data of the scattering potential are equal.
Our results show that in dimensions higher than two, all Fourier coefficients appearing in the equations are, at least in the generic case, uniquely recoverable from the measured data. In two dimensions, however, only a subset of the coefficients is uniquely determined, while on the remaining region distinct values for the Fourier coefficients produce identical measurements.
@@ 7. Conclusion
Nevertheless, even in that case, the recoverable Fourier coefficients go beyond those that can be directly read off from the Fourier data of the measurments.

### v2

@@ Abstract
Diffraction tomography aims to recover an object’s scattering potential from measured wave fields. In the classical setting, the object is illuminated by plane waves from many directions, and the Fourier diffraction theorem provides a direct relation between the Fourier transform of the object’s scattering potential and the Fourier transform of the measurements.
In many practical imaging systems, however, focused beams are used instead of plane waves. These beams are then translated across the object to bring different regions of interest into focus.
This article discusses what information about the scattering potential can be extracted from such measurements. As in the classical case, the analysis is based on a recently derived Fourier diffraction relation that relates the measurements to the Fourier coefficients of the scattering potential. However, this relation does not immediately provide an explicit reconstruction formula, but instead leads to a linear equation system for the Fourier coefficients.
We therefore prove in this work that all Fourier coefficients appearing in these relations are in dimensions higher than two generically uniquely determined. In the two-dimensional case, on the other hand, only a particular subset of the Fourier coverage is uniquely recoverable, while on the remaining region distinct coefficients may produce identical data.
@@ 1. Introduction
Building on this framework, the present article focuses on the mathematical structure of the resulting Fourier diffraction relation. More precisely, we investigate whether, and under which conditions, the Fourier coefficients of the scattering potential can be uniquely recovered from the resulting system of linear equations.
While the physically relevant case is clearly the three-dimensional setting, we discuss the problem in arbitrary dimensions, with particular attention to the theoretically interesting two-dimensional case where the dimensions of the gathered data and the Fourier data of the scattering potential are equal.
Our results show that in dimensions higher than two, all Fourier coefficients appearing in the equations are, at least in the generic case, uniquely recoverable from the measured data. In two dimensions, however, only a subset of the coefficients is uniquely determined, while on the remaining region, distinct values for the Fourier coefficients produce identical measurements.
These findings are essential for the reconstruction process of raster scan diffraction tomography: only those Fourier coefficients that are uniquely determined by the measurement data can be incorporated into backpropagation. The established invertibility conditions therefore provide the mathematical foundation for guaranteeing unique reconstructibility of the scattering potential within the accessible frequency coverage.
@@ 7. Conclusion
Nevertheless, even in that case, the recoverable Fourier coefficients go beyond those that can be directly read off from the Fourier data of the measurements.